OpenStudy (happpily):

Can someone please help me find what is wrong with my answer? I can't figure it out :-(

OpenStudy (happpily):

I had to find the slope of my problem so I did this: 843-431/4-18 which gave me 412/14. I simplified and got 206/7 which equals 29.42. What is wrong with what I did ???

OpenStudy (happpily):

My friend tells me it's my subtraction but I use a calculator and it's still right?

OpenStudy (s4sensitiveandshy):

what's the question ?

OpenStudy (s4sensitiveandshy):

if the fraction is correct then the mistake is *4-18* isn't equal to positive 14

OpenStudy (anonymous):

\[\frac{843-431}{4-18}=-\frac{206}{7} \]

OpenStudy (happpily):

Negative? Why :o

OpenStudy (s4sensitiveandshy):

subtraction: when the bigger number is negative then the answer would be negative 4-9= -5 2-10= -8

OpenStudy (happpily):

but none of them were negatives, all positives

OpenStudy (s4sensitiveandshy):

alright then what's the question ? there is no question.

OpenStudy (happpily):

If I simplified the. It would be -29.42 ?

OpenStudy (s4sensitiveandshy):

maybe you didn't substitute dthe #s right...?

OpenStudy (happpily):

But how did @robtobey get a negative answer and 206/7 from 843-431/4-18?

OpenStudy (s4sensitiveandshy):

-412/14 =- 29.4285 if you want to round it nearest hundredth then it should be -29.43 Because the number after hundredth term is greater than 5 so you should round it.

OpenStudy (s4sensitiveandshy):

he simplified the fraction \[\ \frac{ -206 }{ 7 }=\frac{-412}{14}\] divide the numerator and denominator by 2 you will get 206/7

OpenStudy (happpily):

Okay I understand it, thank you.

OpenStudy (s4sensitiveandshy):

no problem. good luck!

OpenStudy (happpily):

I have to find the line can you please check my answer to that one to.

OpenStudy (s4sensitiveandshy):

sure

OpenStudy (happpily):

Thank you so much again!! Okay here's my equation: 14=-29.42(18)+b 14= -529.56 B= -543.56

OpenStudy (s4sensitiveandshy):

first of all it is 29.43 remember rounding ? \[14=-29.43(18)+b\] multiply by 18 we get \[14=-529.74+b\] now you should add 529.74 to solve for x. how did you get negative answer ?